Christa is 7 years younger than Mark. 5 years hence, the sum of their ages will be 31. How old is each?
Answer by fractalier(6550)

Call their ages NOW, c and m. We have
c = m - 7 and
(c + 5) + (m + 5) = 31
Substitute and get
(m - 7 + 5) + (m + 5) = 31
2m + 3 = 31
2m = 28
m = 14
c = 7